2-49 Refer to Problem 2-48 and now assume the prices for the equipment that Constance wants to replace have not been constant. Use the cost index data for each piece of equipment to update the costs to the price that would be paid today. Develop the overall cost for Constance, again assuming the 15% trade-in allowance for the old equipment. Trade-in value is based on original cost. Cost Index Original When Originally Cost Index Purchased Today Equipment Varnish bath Power scraper Paint booth 154 780 49 900 76
Answer #1

Solution:

Equipment Cost of New Equipment Trade-In Value Net cost
Varnish Bath 4,841 (171/154) = $5,375 3,500 (0.15) 4,850
Power Scraper 291 (900/780) = $336 250 (0.15) 298
Paint Booth 6892 (76/49) = $10,690 3,000 (0.15) 10,240
15,338